# A Comprehensive Assessment Benchmark for Rigorously Evaluating Deep Learning Image Classifiers

### v2

@@ 2.3 Rejection Criteria
In order for classifiers to be evaluated using the proposed method, it is necessary to define a criteria that will be used to determine if a sample is accepted for classification, or rejected. Many different methods for unknown class rejection have been proposed (Vaze et al.,, 2022; Pang et al.,, 2022; Chen et al.,, 2022; Hendrycks and Gimpel,, 2017; Tajwar et al.,, 2021). This article will primarily use maximum softmax probability (MSP; Hendrycks and Gimpel,, 2017), which defines the confidence that a sample is of a known class as the maximum response of the network output after application of the softmax activation function. MSP is used as this is a standard, baseline method, used frequently in the literature on unknown class rejection. Furthermore, a comparison of different rejection methods using the same metric on which the proposed benchmark is based found MSP to be one of the best methods (Zhu et al.,, 2024).
@@
Previous work has shown that relative performance of different unknown class rejection methods vary depending on a number of factors (Szyc et al.,, 2023) including the data-sets Tajwar et al., (2021) and models (Zhu et al.,, 2024) used. Hence, to confirm that the results obtained in this study are not unique to MSP results for three other post-hoc methods of unknown class rejection (those that rely on the responses produced by the network) are also reportedbbbOnly post-hoc methods are considered, as (1) they can be applied to the same network, allowing for a fair comparison and avoiding the need to re-train the network to assess different methods, (2) because post-hoc methods have performance comparable to, or better than, other methods (Yang et al.,, 2022). (1) Maximum logit score (MLS; Vaze et al.,, 2022), which defines the confidence that a sample is of a known class as the maximum response of the network output before any activation function is applied. This was shown to be the best method of unknown class rejection in a recent comparison of state-of-the-art methods (Vojir et al.,, 2023). (2) Energy score (Liu et al.,, 2020), which defines confidence as the negative logarithm of the denominator of the softmax activation function applied the network output layer. (3) GEN (Liu et al.,, 2023) a very recently proposed method that claims to be superior to all three other methods considered here.
@@
Consistent with the method proposed by Zhu et al., (2024), a rejection threshold is set such that a fixed percentage correctly classified clean examples are accepted. In other words, such that the confidence associated with those samples (calculated using either MSP, MLS, Energy score or GEN depending on the experiment being performed), is greater than or equal to the threshold. All samples, from any data-set, with confidence lower than the threshold are rejected, while all others are accepted and the class label predicted.
2.4 Choice of Models to Assess
The aim of the current work is to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods advocated above. However, it is clearly impossible to attempt to evaluate the performance of every network architecture or to evaluate the huge variety of different techniques that has been proposed for improving adversarial robustness, generalisation to common corruptions, or for unknown class rejection.
ResNet18 (He et al.,, 2016) was chosen as the principal DNN architecture to test as this is very commonly used in the literature. This model has just over 11M trainable parameters. Additional experiments were also performed using WideResNet34-20 (Zagoruyko and Komodakis,, 2016) which has just under 193M parameters. For the MNIST data-set smaller networks were used. Firstly, a full-connected multi-layer perceptron network, containing 200 neurons in each of 3 hidden layers (denoted as MLP200-200-200). This network used the ReLU activation function and had 239,410 trainable parameters. Secondly, a fully-convolutional neural network (ConvNet) containing 5 layers, with 32 3x3 convolutional filters in each hidden layer, and average pooling of stride 2 applied after the first and second layers. The ReLU activation function was applied to each layer (except the final output). This network is denoted as ConvNet5-3-32 where the numbers indicate the depth, kernel-size, and width respectively. This network contained 30,954 trainable parameters.
- baseline
Improved performance on the standard test data is often achieved by applying simple geometric data augmentations to the training data. Here 4-pixel random cropping and random horizontal flipping where applied to the CIFAR10, CIFAR100 and TinyImageNet training data, while no augmentations were applied to the MNIST data. These standard augmentations were used in combination with the additional augmentation methods listed below. Given that these baseline augmentation methods are very commonly used when training models, the results show how standardly trained networks cope under comprehensive evaluation. The results for this method of training also provides a baseline against which to judge the effectiveness of other methods that have been proposed to improve generalisation performance.
- noise
Augmenting training images with random noise has been shown to reduce over-fitting, and hence, improve generalisation (Sietsma and Dow,, 1991; Holmstrom and Koistinen,, 1992; Goodfellow et al.,, 2016), especially to common image corruptions (Rusak et al.,, 2020; Gilmer et al.,, 2019). Various methods of adding noise to the training data have been proposed. Rusak et al., (2020) augmented half the training data with Gaussian noise, and found that this technique produced performance on corrupt test data that was competitive with the more complex augmentation method, AugMix, and that for the MNIST data-set specifically was just as effective as their proposed method of Adversarial Noise Training. Lopes et al., (2019) advocated adding Gaussian noise to only a randomly chosen patch of the image, which they claimed produced better robustness to image corruptions on CIFAR10 data than applying random noise to the whole image. Gilmer et al., (2019) used an alternative method for corrupting training images with Gaussian noise in which the standard deviation of the noise applied to each image was chosen with equal probability from the range . The method implemented here is a generalisation for the method proposed by Gilmer et al., (2019). Specifically, for each image the standard deviation of the noise was chosen from a uniform distribution with range . Then each pixel intensity value was independently modified by the addition of a value chosen at random from a Gaussian distribution with mean of zero and the chosen standard deviation. The original images had pixel intensity values in the range [0,1], and after the application of the noise pixel values were clipped at 0 and 1 to remain in this range. Appropriate values of and were found via a grid search using small ConvNets. Using this method and were chosen for use with CIFAR10, and the same hyper-parameters were also used for CIFAR100 and TinyImageNet. For MNIST was used.
- AT
Multi-step adversarial training with Projected Gradient Descent (PGD; Madry et al.,, 2018) is a standard, and highly effective, defence against adversarial attack. As a result, this method has become a standard benchmark against which all other methods of adversarial defence are judged. This method was, therefore, also selected as a data augmentation technique to be tested. It was implemented using the torchattacks PyTorch library (Kim,, 2021). To perform adversarial training, clean images from the current batch of training examples were used to generate non-targeted adversarial examples, that then replaced the original training images (as in Madry et al.,, 2018). Following standard protocols, the maximum allowed perturbation was constrained by the -norm to be less than the perturbation budget (), and the training-time adversarial images were generated using 10 steps of PGD (denoted as PGD for brevity). Standard values used in the previous literature were used for . Specifically, was set to for performing AT with the CIFAR10, CIFAR100 and TinyImageNet training data, and was used for MNIST.
- pixmix
- regmixup
In addition, evaluations were also performed on current state-of-the-art robust pre-trained models available from the RobustBench model zooccchttps://github.com/RobustBench/robustbench (Croce et al.,, 2021).
2.5 Experimental Setup
Training was performed using a batch size of 128. For the CIFAR10, CIFAR100 and TinyImageNet data-sets, training was performed for 110 epochs using the SGD optimiser with weight decay of 5e-4 and an initial learning rate of 0.1. The learning rate was step-wise decayed by a factor of 10 at the start of epochs 100 and 105 for the CIFAR data-sets, and at epochs 45 and 89 for TinyImageNet. The set-up for the CIFAR data-sets was based on that recommended for AT by Pang et al., (2021). Hence, these hyper-parameters may not be optimal for training with the other data augmentations used. For the MNIST data, 20 epochs of training using the Adam optimiser and a constant learning rate of 0.001 was used.
For each experimental condition (combination of data-set, network architecture, and training data augmentation method) the network was trained and evaluated three times, each time with a different random weight initialisation and random presentation order of training samples. The reported result for each condition is the mean accuracy over these three trials. The standard-deviation in the mean performance measured using the proposed metric is shown in the tables of results and generally shows a high degree of consistency across trials. Exceptions to the above, where only a single trial was performed, were the experiments with WideResNet34-20 and the experiments performed using a publicly available checkpoint created by other researchers.
Open-source code written in Python using the Pytorch library (Paszke et al.,, 2019), which performs all the experiments described in this article, is available for download from: https://codeberg.org/mwspratling/RobustnessEvaluation.
@@ 3 Results
3.1 Comparison with Standard Metrics
@@
| Method | Accuracy (%) | AUROC (%) | |||||||||
@@
| Accuracy (%) | AUROC (%) | ||||||||||
[…]
| noise | |||||||||||
[…]
| regmixup | |||||||||||
[…]
| noise | .45) | |||||
[…]
| regmixup | .28) |
[…]
| noise | .09) | |||||
[…]
| regmixup | .97) |
Table 2 shows results for ResNet18 networks trained on CIFAR10 using different training-time data augmentation methods that have been proposed to improve performance (as discuss in Section 2.4). Table 3 shows equivalent results for CIFAR100 trained networks. Evaluation of the performance of these networks using standard metrics is shown in Figs. 3(a) and 3(d). It can be seen that, as expected, AT with PGD produces the best defence against adversarial attacks with AA. It can also be seen that, compared to a network trained with the baseline augmentations, AT causes a reduction in the accuracy with which clean data is categorised and a reduction in performance on unknown class rejection. Also not unexpectedly, pixmix produces the best performance on unknown class rejection for all the data-sets (except CIFAR10 for the CIFAR100 trained network), and produces the best performance on the corrupt data.
These findings are also captured by the new evaluation metric, results for which are shown in Figs. 3(b) and 3(e) when the rejection threshold is set such that 99% of correctly classified clean examples are accepted. The results summarising performance on each type of data show that AT produces the highest accuracy for adversarial samples, and pixmix produces the highest accuracy on corrupt data and the two types of unknown class data (novel and unrecognisable). The trade-offs in performance over different types of data can also be seen. The average performance over all data types (shown in the last columns of each table) suggests that pixmix produces the best performance overall.
The results for all methods on unknown class data is far worse than the standard metrics would imply. This is because the same rejection threshold has been used for all data types, rather than being tuned to optimise performance on unknown class rejection. This illustrates that current metrics for unknown class rejection grossly over-estimate likely real-world performance (as discussed in Section 1.5). In contrast, the performance of all methods on known class data (clean, corrupt and adversarial) is higher than would be expected from the standard metrics. This is because networks can reject samples for which prediction confidence is low which include many samples for which the predicted class is wrong.
Note that these results generally show that current methods of training deep networks produce classifiers that lack robustness. For most of the methods evaluated (which includes methods that are designed to produce state-of-the-art robustness) the resulting network can only make the correct decision on around 50% of the samples for CIFAR10, which is a relatively simple data-set. Even pixmix, that performs best overall, only makes accurate decisions for just over 71% of the test data. Furthermore, it might reasonably be argued that a model is only as robust as the worst results would suggest, as a malicious actor can select the type of data used to attack a system, and in the real world a model might naturally encounter all types of data and would need to deal correctly with all. For each method of training investigated here there is at least one type of test data on which the model performs pitifully. For example, AT can only correctly reject 2.2% of unrecognisable images: all the others are given incorrect class labels with a high confidence. Massive amounts of time and resources have gone into developing techniques to enhance adversarial robustness. The current results suggest that this effort may have been wasted, as an attacker can easily and effectively induce a classifier to generate the wrong predictions by presenting it with easily generated synthetic images. The results in Section 3.5 show that the most recent state-of-the-art methods of adversarial robustness, while having better performance on unknown class rejection, are still far from robust to such data.
Figs. 3(c) and 3(f) show results for the proposed evaluation method when the rejection threshold is set such that 95% of correctly classified clean examples are accepted (an increased rejection threshold compared to previously). For the CIFAR10 trained networks, rejecting more of the clean samples results in a reduction in performance on the clean data as some samples that were correctly classified have been rejected. In contrast, the results for CIFAR100 show that performance on the clean data improves with the increased rejection threshold, due to the higher proportion of clean samples that are mis-classified by the CIFAR100 trained networks. For both data-sets, rejecting more samples for which confidence in the prediction is low results in a large improvement in performance on the adversarial, novel and unrecognisable data types for all training methods.
The results produced with both thresholds are consistent showing similar trade-offs in performance on the different data types for each of the tested methods, and producing similar rankings of different training methods, although the differences between methods is somewhat reduced when the threshold is set so as to accept 95% of correctly classified clean samples. To avoid redundancy, and for brevity, only results for one threshold are reported subsequently. The 95% threshold was chosen in order to present the most favourable performance for the tested DNNs.
3.2 Different Unknown Class Rejection Methods
@@
The same networks evaluated in the previous section were re-evaluated using MLS, Energy score and GEN as the criteria for rejection (see Section 2.3). The results for the networks trained on CIFAR10 are shown in Table 4, and for CIFAR100 in Table 5. Comparing the results in Table 4 with the results obtained using MSP shown in Fig. 3(c) it can be seen that while individual accuracy values can vary, a similar overall performance is produced using each method of rejection, and as a result a similar ranking of the overall performance of each training method is obtained. The same can be said for the CIFAR100 results shown in Table 5 and Fig. 3(f). The results obtained, therefore, do not seem to be specific to the particular method of rejection used, and to avoid redundancy, and for brevity, only results for MSP are reported subsequently. This choice was made as the worse performance on any data type is better for MSP than for any of the other three methods, and hence, MSP seems to have a slight advantage in improving robustness.
| Training Method | Clean | Corrupt | Adversarial | Novel | Unrecog. | Mean |

### v3

@@ 2.3 Rejection Criteria
In order for classifiers to be evaluated using the proposed method, it is necessary to define a criteria that will be used to determine if a sample is accepted for classification, or rejected. Many different methods for unknown class rejection have been proposed (Vaze et al.,, 2022; Pang et al.,, 2022; Chen et al.,, 2022; Hendrycks and Gimpel,, 2017; Tajwar et al.,, 2021). The proposed benchmark uses maximum softmax probability (MSP; Hendrycks and Gimpel,, 2017), which defines the confidence that a sample is of a known class as the maximum response of the network output after application of the softmax activation function. MSP is used as this is a standard, baseline method, used frequently in the literature on unknown class rejection. Furthermore, a comparison of different rejection methods using the same metric on which the proposed benchmark is based found MSP to be one of the best methods (Zhu et al.,, 2024).
@@
Previous work has shown that relative performance of different unknown class rejection methods vary depending on a number of factors (Szyc et al.,, 2023) including the data-sets (Tajwar et al.,, 2021) and models (Zhu et al.,, 2024) used. Hence, to confirm that MSP is an appropriate choice, results for three other post-hoc methods of unknown class rejection (those that rely on the responses produced by the network) are also reported in Section 3.1.1: (1) Maximum logit score (MLS; Vaze et al.,, 2022), which defines the confidence that a sample is of a known class as the maximum response of the network output before any activation function is applied. This was shown to be the best method of unknown class rejection in a recent comparison of state-of-the-art methods (Vojir et al.,, 2023). (2) Energy score (Liu et al.,, 2020), which defines confidence as the negative logarithm of the denominator of the softmax activation function applied the network output layer. (3) GEN (Liu et al.,, 2023) a recently proposed method that claims to be superior to all three other methods considered here. Note, only post-hoc methods are considered, as (A) they can be applied to the same network, allowing for a fair comparison and avoiding the need to re-train the network to assess different methods, (B) because post-hoc methods have performance comparable to, or better than, other methods (Yang et al.,, 2022).
@@
Consistent with the method proposed by Zhu et al., (2024), a rejection threshold is set such that a fixed percentage correctly classified clean examples are accepted. In other words, such that the confidence associated with those samples, is greater than or equal to the threshold. All samples, from any data-set, with confidence lower than the threshold are rejected, while all others are accepted and the class label predicted. Zhu et al., (2024) propose setting the threshold such that either 95% or 99% of correctly classified clean examples are accepted. For the proposed benchmark 95% is used, but results for the alternative are also reported in Section 3.1.2. Instead of using a fixed threshold, it would also be possible to calculate performance over a range of thresholds, as with the AUROC and AUPR metrics. However, in a practical application it would be necessary to define a fixed threshold, so the proposed approach is more likely to accurately reflect real-world performance. Furthermore, the creator of the classifier is likely to want it to perform well on the data it was trained on, so will want to accept a high proportion of the correctly classified clean samples.
@@ 3 Results
Section 3.1 reports results of experiments designed to validate the proposed benchmark and justify some of the choices that have been made in defining it (Sections 3.1.1 and 3.1.2). Section 3.2 reports results designed to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods being advocated. The proposed benchmark can be used to assess the performance of any classifier trained on any of the supported data-sets. Obviously, it would be impossible in this one publication to attempt to benchmark the performance of every such classifier or to evaluate the huge variety of different techniques that can be used with such classifiers to improve adversarial robustness, generalisation performance, or unknown classes rejection. Hence, results are only reported for a few, select, models.
The first set of experimental results (Section 3.1) were produced using self-trained ResNet18 models (He et al.,, 2016). These models were trained on CIFAR10 and CIFAR100 using a variety of training-time data augmentations that have previously been proposed to improve robustness or generalisation. Augmentation methods that produce different levels of robustness to different test data types were chosen, so as to be able to ensure that this variation in performance is captured by the proposed benchmark. These networks have been trained using standard procedures full details of which are given in Appendix A.
The second set of results (Section 3.2) focus of benchmarking the performance of state-of-the-art (SOTA) robust models. These models were taken from the RobustBench leaderboards. RobustBench (Croce et al.,, 2021) is an existing benchmark for assessing the robustness of DNNs. Unlike the proposed benchmark, RobustBench evaluates performance of a model using only a single type of data: either adversarial (produced using either or constrained AA) or common corruptions. From each of these RobustBench leaderboards a selection of the top performing models were evaluated using the proposed benchmark. The selected models included the highest ranked model from each leaderboard for which pre-trained weights were available from the RobustBench repositoryccchttps://github.com/RobustBench/robustbench on November 15th 2024. Other tested models were selected from the top-10 with a bias towards the highest ranked models, but lower ranked models were sometimes selected to increase the diversity of network architectures and robustification methods evaluated.
3.1 Experiments to Validate the Benchmark
This paper advocates for (1) more comprehensive evaluation of DNN performance (Section 1.4), and (2) using a single, consistent, metric to assess performance on all types of testing data (Section 1.5). Figs. 3(a), 3(a), 3(a) and 3(e) shows performance for many different test data sets, measured using existing metrics. Hence, these results address aim (1) but not aim (2). As discussed in Section 1.4 such comprehensive assessment is rarely performed. However, it provides valuable insights into both the strengths and weaknesses of different classifiers. For example, it can be seen that the networks trained using adversarial training (AT) with PGD are most resistant to adversarial attacks with AA, but that such networks have reduced performance on unknown class rejection (and are particularly vulnerable to making incorrect classifications when presented with images containing scrambled pixels). The former is expected, but the latter is ignored in the adversarial robustness literature as evaluations with such data are never performed. Similarly, it is to be expected given previous research that pixmix produces excellent performance on unknown class rejection, however, the vulnerability of such networks to adversarial attack has been mostly undocumented.
@@ 3.1 Experiments to Validate the Benchmark
| Method | Accuracy (%) | AUROC (%) when MSP used as prediction confidence score | |||||||||
@@
| baseline | .77) | |||||
| noise | .99) | |||||
[…]
| Accuracy (%) | AUROC (%) when MSP used as prediction confidence score | ||||||||||
[…]
| noise | |||||||||||
[…]
| regmixup | |||||||||||
[…]
| noise | .09) | |||||
[…]
| regmixup | .97) |
@@
[…]
| noise | .45) | |||||
@@
| regmixup | .28) |
Figs. 3(b) and 3(f) show the assessment of the same networks produced using the proposed benchmark. Hence, these results address both concerns (1) and (2) described in Sections 1.4 and 1.5. The new evaluation metric successfully captures the same pattern of relative performance that can be seen when using the standard metrics. For example, AT produces the highest DAR for adversarial samples, but this improvement comes at the cost of reduced performance (compared to the baseline) in clean accuracy and novel and unrecognisable image rejection. Of the tested augmentation methods, pixmix produces the highest DAR on corrupt data and the two types of unknown class data, which is consistent with the results produced using the standard metrics. regmixup produces the highest performance on clean test data, a result which is captured by both the proposed and standard metrics.
The proposed metrics differ from the standard ones in some respects. For example, for all the tested data augmentation techniques, performance on unknown class data is far worse when measured using DAR than the standard metrics would imply. This is because the same rejection threshold has been used for all data types, rather than being tuned to optimise performance on unknown class rejection. This illustrates that current metrics for unknown class rejection grossly over-estimate likely real-world performance (as discussed in Section 1.5). In contrast, the performance of all the tested models on known class data (clean, corrupt and adversarial) is generally higher than would be expected from the standard metrics. This is because networks can reject samples for which prediction confidence is low which include many samples for which the predicted class is wrong.
An advantage of the proposed benchmark is the use of consistent units to measure performance on each type of test data. This facilitates the comparison of performance of different data types. Using a consistent performance metric for each test data type also allows an average to be calculated. In contrast, it would be meaningless to average the standard metrics given that they are reported in different units. The average performance over all data types (shown in the last columns of Figs. 3(b) and 3(f)) suggests that, of the tested data augmentation methods, pixmix produces the best performance overall. While the average performance helps to identify the most promising methods of improving robustness, the individual results for each test data type help identify remaining vulnerabilities that it would most useful to address in future research.
3.1.1 Choice of Unknown Class Rejection Method
The proposed benchmark uses MSP as the confidence score used to decided if samples are accepted or rejected. The same networks evaluated in the previous section were re-evaluated using alternative rejection methods: MLS, Energy score and GEN (see Section 2.3). The results for the networks trained on CIFAR10 are shown in Fig. 3(c), and for CIFAR100 in Fig. 3(g). Comparing the results in Fig. 3(c) with the results obtained using MSP shown in Fig. 3(b) it can be seen that while individual accuracy values can vary, a similar overall performance is produced using each method of rejection, and as a result a similar ranking of the overall performance of each training method is obtained. The same can be said for the CIFAR100 results shown in Fig. 3(g) and Fig. 3(f). The benchmark’s results, therefore, do not seem to be specific to the particular method of rejection used. Furthermore, the choice of MSP can be justified as the worse performance on any data type is better for MSP than for any of the other three methods, and hence, MSP seems to have a slight advantage in improving robustness.
3.1.2 Choice of Unknown Class Rejection Threshold
The proposed benchmark sets the rejection threshold such that 95% of correctly classified clean examples are accepted (see Section 2.3). Figs. 3(d) and 3(h) shows results for the same networks tested earlier but using a threshold set so that 99% of correctly classified clean examples are accepted. The results produced with both thresholds are consistent showing similar trade-offs in performance on the different test data types, and producing similar rankings of different training methods. The benchmark’s results, therefore, do not seem to be specific to the particular rejection threshold used.
In general setting the threshold so that 95% of correctly classified clean examples are accepted produces more favourable evaluation of each classifier. This is to be expected, as lowering the percentage of correctly classified clean examples that are accepted, increases the rejection threshold and allows more samples associated with low confidence predictions to be rejected. Such low confidence prediction are generally more likely to be wrong, so rejecting them is beneficial.
3.2 Experiments to Evaluate Model Robustness
The results in this section aim to determine how well existing DNNs perform when evaluated using the more comprehensive assessment methods advocated in this paper. Classifiers with SOTA robustness, as determined by RobustBench (Croce et al.,, 2021), are evaluated and the results are shown in Table 5. The comprehensive assessment provides insights into these models that are not unavailable with the narrower analysis provided by RobustBench. For example, the CIFAR10 and CIFAR100 models that have been trained to produce good performance on the common corruptions data have surprisingly good performance at unknown class rejection (Figs. 3(i) and 3(j), upper rows). Hence, despite their relative susceptibility to adversarial attack, their overall performance is competitive with CIFAR10/100 models that have been trained for adversarial robustness.
The CIFAR10 and CIFAR100 models ranked highly by RobustBench for robustness to constrained adversarial attacks are generally very poor at rejecting unrecognisable images (Figs. 3(i) and 3(j), lower rows). As a specific example, RobustBench’s top performing model on CIFAR100 adversarial robustness (Wang et al.,, 2023) produces the wrong predictions around 94% of the time on such images. Hence, these models lack the robustness that conventional methods of evaluation suggest they have. On the contrary, the proposed comprehensive evaluation shows that SOTA methods of improving performance against adversarial attack fail to provide security as an attacker can use cheaply and easily generated random images to fool the classifier with a high chance of success.
The CIFAR10 models developed to have robustness to constrained adversarial attacks, compared to their counterparts trained for robustness, are much better at dealing correctly with samples from unknown classes (Fig. 3(i), middle rows). This, perhaps, suggests that diversifying the methods used to produce training-time adversarial images might help reduce this particular vulnerability. Another approach to improving performance is also suggested by the performance of the XCiT-L12 model (Debenedetti et al.,, 2023) (Fig. 3(j), last row). This is the only transformer architecture to make it into the top-10 of any of the CIFAR10/100 RobustBench leaderboards (at ninth place for CIFAR100 adversarial robustness). This model has far better performance at rejecting unrecognisable images, suggesting that architectural changes may be helpful. However, this improvement in robustness to unrecognisable images comes at the cost of increased vulnerability to novel images. As a result, an attacker could present this model with random images of objects with classes that differ from those in the training data-set and the model would fail 66% of the time. Overall, the proposed comprehensive assessment benchmarks ranks this model as the most robust of the CIFAR100 models tested: a very different ranking than that produced by the narrow testing performed in RobustBench.
It is remarkable how little improvement in overall, average, performance is shown by the SOTA CIFAR10 and CIFAR100 models (Figs. 3(i) and 3(j)) compared with the corresponding results in Figs. 3(b) and 3(f) for (generally) much simpler models made robust with (generally) older methods. Indeed, the best performing CIFAR100 RobustBench model has worse performance, on average, than the best performing model in Fig. 3(f). This is due to the SOTA models failing to make progress on improving performance on all the different types of test data. For example, if we compare the models trained to be adversarially robust (bottom three rows of Fig. 3(j) and the row labelled AT in Fig. 3(f)) it can be seen that the RobustBench models have substantially improved performance on adversarial and clean images, but (as noted above) they remain poor at correctly rejecting samples from unknown classes. Massive amounts of time and resources have gone into developing techniques to enhance adversarial robustness beyond that produced by AT with PGD (Madry et al.,, 2018). The current results suggest that much of this effort may have been wasted, as these models remain vulnerable as an attacker can easily and effectively induce a classifier to generate the wrong predictions by presenting images from unknown classes.
The performance of ImageNet1k trained models (Fig. 3(k)) compared with that of the CIFAR10 models (Fig. 3(i)) suggests that the benefits of an increased training data-set size is off-set by the increase in task difficulty, so that the performance of the ImageNet1k models is no better than that of the CIFAR10 ones. The ImageNet1k results, and all the others that have been presented here, show that current methods of training deep networks produce classifiers that lack robustness. In the real world a model might naturally encounter all the diverse types of data that are used in the proposed comprehensive benchmark and would need to deal correctly with all of them. However, the best performing models would only make accurate decisions for just over 77% of such data. Furthermore, for each of the models evaluated here there is at least one type of test data on which the model performs poorly, which could be exploited by a malicious actor to reduce performance even further.
| Model | Params. | Robustness | Clean | Corrupt | Adv. | Novel | Unrecog. | Mean |